How many litres of water must be added to 10 litres of 30% alcohol in order to produce 50litres of 10% alcohol?
Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52781)

.
Unresolvable problem.

There is NO solution.

Mission is impossible.

Sorry.


Answer by Alan3354(69443)

How many litres of water must be added to 10 litres of 30% alcohol in order to produce 50litres of 10% alcohol?
----------------
You add 40 liters to get 50 liters.
10*30 + 40*0 = 50C
Concentration = 6% tho.
------------
To get 10% :
10*30 + L*0 = (10+L)*10
300 = 10L + 100
L = 20
--> 30 liters of 10%